22west:

Evaluate the expression. ​ 38+16⋅12÷2−(30⋅2) ​

SmokeyBrown:

Ok so parentheses go first, (30*2) = a Then we'll do multiplication and division 16*12/2 = b And last addition and subtraction 38 + b - a And that should give you the answer

22west:

30*2=60

22west:

Am I aloud to use a calculator?

SmokeyBrown:

I don't see why not

22west:

16*12=192

22west:

192/2=96

SmokeyBrown:

yup looks good

22west:

So now 38+60-96?

SmokeyBrown:

That's right

22west:

38+60 is 98 right?

SmokeyBrown:

Yup

22west:

98+96=194?

SmokeyBrown:

That's true, but don't forget we're subtracting 96 from 98 So your equation should be 98 - 96

22west:

Oh subtracting? Then that would be 2

SmokeyBrown:

There we go. Good work!